# Generative AI-Driven Human Digital Twin in IoT-Healthcare: A Comprehensive Survey

### v1

@@ I-A Background
The advancements of Internet of the Things (IoT) in recent years is leading to a paradigm shift in the healthcare industry, which are termed IoT-healthcare [1, 2]. An Ireland healthcare organization estimates that roughly 17.5 million lives are lost annually due to inefficiencies in health data [3]. The main inefficiency is the delay access to and analyze health data, resulting in untimely interventions. Fortunately, the progression of IoT-healthcare presents a promising avenue to tackle these challenges. By enabling the real-time collection and transmission of pertinent health data to servers for in-depth healthcare analysis via personal IoT devices, it can prompt timely healthcare alerts and proactive interventions for saving lives. It has been widely applied in healthcare monitoring, including blood glucose monitoring, cardiac monitoring, respiration monitoring and blood pressure monitoring [4, 5].
@@
Human digital twin (HDT), a promising technology and game changer for IoT-healthcare, is taking this field to another level. HDT can create a digital replica of a human body, comprehensively and precisely characterizing each individual in the digital space while reflecting its physical status in real time [6, 7, 8, 9, 10]. Besides, with visualization and interaction characteristics, HDT is envisioned as a versatile and vivid human digital testbed for revolutionizing the IoT-healthcare, beyond healthcare monitoring applications mentioned above. For instance, a patient’s HDT can be implemented in silico treatment simulations and experiments, facilitating the development of finely-tailored, personalized treatment plans [11]. Additionally, a doctor’s HDT with expert-level medical knowledge can be a personal 24/7 doctor to answer the patients’ queries [12].
@@
The successful establishment and implementation of HDT largely depend on the high-fidelity human modeling, supported by comprehensive individual-level data encompassing appearance, movement, and physiological data, acquired from multi-source, such as IoT devices. In addition, as an intelligent human digital testbed, HDT needs to generate various human-like feedback during immersive real-virtual interactions. These include, for example, providing intuitive feedback on drug-disease responses, and simulating haptic feedback to replicate the tactile sensations experienced by humans in real-world scenarios. All such requirements are difficult to meet due to several crucial reasons, e.g., data scarcity, bias, noise and intricate digital modeling. Fortunately, generative artificial intelligence (GAI) has been recognized as a promising technology that can effectively fulfill or assist the implementation of HDT for IoT-healthcare [13]. GAI can leverage advanced AI algorithms to automatically create, manipulate, and modify valuable while diverse data [14, 15]. Specifically, GAI models, such as generative adversarial network (GAN), variational autoencoder (VAE), transformer, and diffusion model, with their powerful creativities and data analysis abilities can generate ultra-realistic individual-level data and make informed decisions for HDT in IoT-healthcare, which will be elaborated in Section III. Thus, our survey focuses on how GAI enables HDT in IoT-healthcare, namely GAI-driven HDT in IoT-healthcare.
I-B Related Work and Contributions
@@
Given the increasing interest of GAI-driven HDT in IoT-healthcare, several surveys and tutorials have been recently published [16, 17, 18, 19, 20, 21]. Table I presents a comparison of these related works compared with ours.
TABLE I: Comparison of The Related Work with Our Survey.
@@
The surveys above are centered around IoT-healthcare, HDT and GAI. However, none of them explore the potential of GAI with remarkable data generation and analysis capabilities for enabling HDT in IoT-healthcare applications. This motivates us to compose this survey investigating the GAI-driven HDT in IoT applications. The contributions of this survey can be summarized as follows:
@@
The rest of this paper is organized as follows: Section II presents the fundamentals of HDT and GAI, and gives an overview of the framework of GAI-driven HDT. In Section III, the implementation of GAI-driven HDT is analyzed in detail. Section IV surveyed the IoT-healthcare application of GAI-driven HDT, including personalized health monitoring and diagnosis, personalized prescription, and personalized rehabilitation. Section V explores several future research directions of GAI-driven HDT in IoT-healthcare. Section VI concludes this survey paper.
@@ II-A Human Digital Twin
HDT, focuses on digital replicas of human beings while the conventional DT limits the attention to non-living physical entities, e.g., machines [27] and networks [28]. Then, several distinguishing characteristics between HDT and the conventional DT are detailed below and summarized in Table II.
TABLE II: Differences Between HDT and the Conventional DT.
@@
  • Mobility pattern: Unlike the position-fixed machines in the conventional DT, the mobility patterns in HDT may be highly predictable. The mobility patterns of human beings can be categorized into human positional and postural mobility. Positional mobility, like a person moving from indoors to outdoors, may cause radio frequency (RF) propagation characteristics to change and even the service migrations. Additionally, the postural mobility of a human, like lying, sitting, walking, may cause signal strength to fluctuate due to the influence of human bodies on the path loss, known as body shadowing [8, 9, 13].
@@ II-B GAI Techniques
In this subsection, we delve into the recent primary trends of GAI models, including generative adversarial network, variational autoencoder, transformer and diffusion model [14, 35]. These models find frequent applications in HDT for IoT-healthcare.
@@
GAN has been successfully applied in HDT for IoT-healthcare. For instance, a heart DT model was integrated into optimizing a GAN to generate electrocardiogram (ECG) data [37], which can be used to solve the ECG data scarcity issue when digital modeling of cardiac activities. Additionally, a chained GAN-based approach, which connected multiple GAN models, was proposed to simulate the pathology of tissues in the human body. This approach can be used to model tissue digital twin in HDT, as well as help diagnose diseases and predict their progress [38].
@@
VAE has been successfully applied in HDT for IoT-healthcare. For instance, a VAE-based approach was introduced to reconstruct complete body movements using signals from the PT’s head-mounted device (HMD) [40]. This method enabled the digital reconstruction of full-body motion based on signals from the head and hands. It proved beneficial for accurately modeling human motions in scenarios where motion data obtained from the PT may inadequately represent complete body movements, thereby aiding motion monitoring endeavors. Moreover, a VAE-based method was proposed to predict potential subsequent steps in the clinical measurement trajectory of a patient encountering an ischemic stroke [41]. This approach holds promise in simulating disease progression within HDT, thereby assisting in customizing treatment plans tailored to individual patients.
@@
Transformer has been successfully applied in HDT for IoT-healthcare. For example, the state-of-the-art (SOTA) language generation model, generative pre-trained transformer (GPT), has been incorporated into electronic health records (EHR) workflows to autonomously respond to patients’ healthcare inquiries [43]. This model can be visualized as an HDT representing a doctor, equipped with extensive medical knowledge, capable of addressing patient queries. Additionally, a pre-trained transformer-based approach has been proposed to concurrently learn gene and cell embeddings, enabling the capture of intricate gene-to-gene interactions at the single-cell level [44]. This approach holds promise in digitally modeling diverse facets of cellular processes within HDT, offering insights into personalized responses to treatments.
@@
Diffusion model has been successfully applied in HDT for IoT-healthcare [14, 46, 47]. For instance, a diffusion mode-based approach has been proposed to generate individual electroencephalogram (EEG) data [46]. It can help solve the EEG data scarcity issue in HDT, and enable HDT to digitally model brain activity patterns, supporting the neurological health monitoring. Additionally, by collaborating with an assistive modality embedding as prior information to diffusion model formulation, a diffusion model-based approach has been proposed for positron emission tomography (PET) denoising [47]. It can be used to denoise the acquired data with noise, providing cleaned data to HDT for better personalized healthcare monitoring and diagnostic accuracy.
@@ II-C Framework of GAI-driven HDT
Data acquisition component is significantly crucial for HDT, which is the fuel of HDT. The required substantial health data, such as EEG, ECG, and magnetic resonance imaging (MRI) data, for HDT commonly collected from the IoT-enabled pervasive sensing and medical institutions [50]. However, these traditional data acquisitions methods are usually inefficient due to various factors. To this end, GAI can generate ultra-realistic health data based on the collected data for enriching the datasets.
The communication component plays a bridge role in HDT, which is responsible for bi-directional data transmissions between the physical and digital worlds, such as the transmission of the collected data and the feedback in the informal world. However, these data are usually large-scale, multi-modal, and time-sensitive, which are hard to fully met by traditional communication systems. To this end, GAI-enabled communication, such as GAI-enabled semantic communication [51, 52, 53] and cross-modal communication [54], can be applied to support communication in HDT by generating the transmitted data at the received sides, enhancing the data transmission performances.
@@
In the following sections, we will delve into implementing GAI-driven HDT in IoT-healthcare.
@@ III GAI-driven HDT Implementation
TABLE III: Summary of GAI-based Approach for Data Acquisition in HDT.
@@ III-A GAI-enabled Data Acquisition
Data forms the cornerstone of HDT development and service provision [8]. Commonly employed data acquisition methods for HDT primarily stem from diverse sources, including medical institutions [55], as well as both non-invasive and invasive sensors equipped in PTs [34]. Nevertheless, these methods present challenges due to their time-consuming, costly, intrusive nature, and limited scalability. This restricts acquiring extensive individual-level data essential for robust HDT development and comprehensive service delivery. GAI can significantly assist the data acquisition in HDT, by offering diverse and highly realistic synthetic data. In the following, we introduce several common synthetic HDT data generated by GAI, including synthetic physiological, medical imaging and motion data, as summarized in Table III.
[…]
@@
The medical imaging data hold significant importance for HDT. Medical imaging provides detailed insights into a PT’s anatomy, allowing for a precise digital representation. This information aids in creating accurate simulations of physiological structures and functions within the VT. Besides, by integrating this data, the VT can simulate and predict the progression of diseases. However, existing medical imaging acquisition methods rely on medical imaging devices, such as MRI or computed tomography (CT) scanners, which encounter high costs, patient discomfort, and safety concerns. To this end, GAI can generate synthetic medical images, augmenting existing datasets without additional patient scans. For instance, Moghadam et al. in [61] proposed a diffusion model-based approach for the synthesis of histopathology images, as shown in Fig. 4. Specifically, the authors used color normalization to force the diffusion model-based approach to learn morphological patterns, and used perception prioritized weighting, aiming to prioritize focusing on diffusion stages with more important structural histopathology contents. Experimental results showed that the proposed approach outperformed the GAN-based approach proposed in [62] by generating high quality histopathology images of brain cancer. Similarly, Yoon et al. in [63] proposed a sequence-aware diffusion model (SADM) for the generation of longitudinal medical images. SADM used a sequence-aware transformer as the conditional module in the diffusion model. This enabled learning longitudinal dependency even with missing data during training and allowed autoregressive generation of a sequence of medical images during inference. Experimental results showed that SADM can generate high quality longitudinal cardiac and brain MRI. While GAN and diffusion model made remarkable progress in medical image generation, they cannot explicitly learn the probability density function of the input data and are highly sensitive to the hyperparameter selections. To mitigate these issues, Hajij et al. in [49] investigated normalizing flows (NFs) based approach as an alternative for synthesizing medical images. Particularly, the authors trained a RealNVP [64], a popular NF model for medical image synthesis, on two medical imaging datasets: chest X-ray [65] and skin cancer [66]. The experimental results showed that the NF-based medical image synthesis approach is an attractive alternative to GAN-based and diffusion model-based approaches.
@@
In summary, Generative AI offers promising solutions to address the limitations of current data acquisition methods in HDT by generating diverse and high-realistic datasets.

### v2

@@ I-A Background
The Internet of the Things (IoT) refers to a group of Internet-connected devices that have computing capability, are easy to identify, and are able to share data over the Internet without the need for direct touch [1]. The IoT progression has led to unmatched connectivity among devices, paving the way for the creation of intelligent environments, enabling new ways of working, communicating, interacting, entertaining, and living [2]. Notably, IoT is leading to a paradigm shift in the healthcare industry, which are termed IoT-healthcare [3, 4]. A global healthcare organization estimates that roughly 17.5 million lives are lost annually due to inefficiencies in collecting and analyzing health data [5]. Fortunately, the advancement of IoT-healthcare presents a promising avenue to tackle these challenges [6]. By enabling the real-time collection and transmission of pertinent health data to servers for in-depth healthcare analysis via personal IoT devices, it can prompt timely healthcare alerts and proactive interventions for saving lives [7, 8]. IoT-healthcare has been widely applied in healthcare monitoring, including blood glucose monitoring, cardiac monitoring, respiration monitoring and blood pressure monitoring, providing people with pervasive healthcare services [6, 7].
Meanwhile, digital twin (DT) has attracted substantial attentions since 2002 when Michael Grieves delivered a presentation at the University of Michigan [9]. With the continuous progression, DT, nowadays, is characterized as a smart, evolving system that precisely mirrors a physical entity at various granularity levels, and monitors, controls, and optimizes the physical objects during its lifecycle [10]. When adopting DT in the human-centric systems, a revolutionary technology known as Human Digital Twin (HDT) has emerged. HDT is a promising technology and game changer for IoT-healthcare, taking this field to another level. HDT can create a digital replica of a human body, comprehensively and precisely characterizing each individual in the digital space while reflecting its physical status in real time [11, 12, 13, 14, 15].
@@
Besides, with visualization and interaction characteristics, HDT is envisioned as a versatile and vivid human digital testbed for revolutionizing the IoT-healthcare, beyond healthcare monitoring applications mentioned above. For instance, a patient’s HDT can be implemented in silico treatment simulations and experiments, facilitating the development of finely-tailored, personalized treatment plans [16]. Additionally, a doctor’s HDT with expert-level medical knowledge can be a personal 24/7 doctor to answer the patients’ queries [17].
@@
The successful establishment and implementation of HDT largely depend on the high-fidelity human modeling, supported by comprehensive individual-level data encompassing appearance, movement, and physiological data, acquired from multi-source, such as IoT devices. In addition, as an intelligent human digital testbed, HDT needs to generate various human-like feedback during immersive real-virtual interactions. These include, for example, providing intuitive feedback on drug-disease responses, and simulating haptic feedback to replicate the tactile sensations experienced by humans in real-world scenarios. All such requirements are difficult to meet due to several crucial reasons, e.g., data scarcity, bias, noise and intricate digital modeling. Fortunately, generative artificial intelligence (GAI) has been recognized as a promising technology that can effectively fulfill or assist the implementation of HDT for IoT-healthcare [18].
GAI can leverage advanced AI algorithms to automatically create, manipulate, and modify valuable while diverse data [19, 20]. Specifically, GAI models, such as generative adversarial network (GAN), variational autoencoder (VAE), transformer, and diffusion model, with their powerful creativities and data analysis abilities can generate ultra-realistic individual-level data and make informed decisions for HDT in IoT-healthcare, which will be elaborated in Section III. Thus, our survey focuses on how GAI enables HDT in IoT-healthcare, namely GAI-driven HDT in IoT-healthcare.
I-B Review Methodology
We identify three specific research questions that lead to the accomplishment of this work in the following.
1) RQ1: What is the definition of HDT? And what GAI models can be potentially applied in HDT?
2) RQ2: How GAI can enable the implementation of HDT?
3) RQ3: How GAI-driven HDT empowers IoT-healthcare?
We perform this survey in a systematic way following the guidelines, preferred reporting items for systematic reviews and meta-analyses (PRISMA) [21], which consists of three phases, i.e., identification, screening, eligibility and included.
1. Identification: For the literature review, we used three scientific platforms: Google Scholar, Scopus and Web of Science. The search string is defined as shown in Table I. Only papers written in English were included. Using these criteria, we found 2,416 papers on Google Scholar, 900 on Scopus, and 1,679 on Web of Science. We then eliminated duplicates from all the papers found. Additionally, some papers that were not highly relevant, potentially due to certain platforms ignoring logical operators (i.e., AND and OR), were also filtered out. The total number of selected papers in this phase was 1,736.
2. Screening: In this phase, we evaluated the selected papers in identification phase on the basis of their titles and abstracts. In this survey, we focus on the application of GAI techniques in building HDT from the perspective of IoT-healthcare. Specifically, GAI techniques are expected to enable each component of HDT, including data acquisition, communication, data management, digital modeling, and data analysis, which will be elaborated in Section III. All of these GAI-driven components are required to potentially empower IoT-healthcare. Thus, following these criteria, we further filtered these 1736 papers to be 369 papers by examining their titles and abstracts.
3. Eligibility and Included: Since the titles and abstracts of some papers were not clear enough to be directly evaluated, a full-text screening was performed on these 369 papers during this phase. We followed the criteria from the screening phase. Finally, a total of 172 papers were included in this paper.
TABLE I: Search String of Acquiring the Literature.
I-C Related Work and Contributions
@@
Given the increasing interest of GAI-driven HDT in IoT-healthcare, several surveys and tutorials have been recently published [22, 23, 24, 25, 26, 27]. Table II presents a comparison of these related works compared with ours.
TABLE II: Comparison of The Related Work with Our Survey.
@@
Unlike the above surveys that broadly cover IoT-healthcare, HDT, and GAI, ours specifically investigates the pivotal role of GAI in HDT within IoT-healthcare by providing comprehensive and up-to-date review of the current works. GAI’s exceptional capabilities in data generation and analysis are crucial for improving all aspects of HDT implementation, from data collection and management to digital modeling and analysis. Our work marks the first in-depth survey of GAI-driven HDT in IoT-healthcare, offering insights into its application in personalized health monitoring adiagnosis, prescription, and rehabilitation. We also highlight current challenges and envisage future directions, ensuring our survey captures the emerging trends in this evolving field. The contributions of this survey can be summarized as follows, which also aims to answer to research questions listed in Section I-B :
@@
The rest of this paper is organized as follows: Section II aims to answer RQ1 in Section I-B, which presents the fundamentals of HDT and GAI, and gives an overview of the framework of GAI-driven HDT. Section III aims to answer RQ2 in Section I-B, where the implementation of GAI-driven HDT is analyzed in detail. Section IV aims to answer RQ3 in Section I-B, which surveys the IoT-healthcare application of GAI-driven HDT, including personalized health monitoring and diagnosis, personalized prescription, and personalized rehabilitation. Section V explores several future research directions of GAI-driven HDT in IoT-healthcare. Section VI concludes this survey paper.
@@ II-A Human Digital Twin
HDT, focuses on digital replicas of human beings while the conventional DT limits the attention to non-living physical entities, e.g., machines [33] and networks [34]. Then, several distinguishing characteristics between HDT and the conventional DT are detailed below and summarized in Table III.
TABLE III: Differences Between HDT and the Conventional DT.
@@
  • Mobility pattern: Unlike the position-fixed machines in the conventional DT, the mobility patterns in HDT may be highly predictable. The mobility patterns of human beings can be categorized into human positional and postural mobility. Positional mobility, like a person moving from indoors to outdoors, may cause radio frequency (RF) propagation characteristics to change and even the service migrations. Additionally, the postural mobility of a human, like lying, sitting, walking, may cause signal strength to fluctuate due to the influence of human bodies on the path loss, known as body shadowing [13].
@@ II-B GAI Techniques
In this subsection, we delve into the recent primary trends of GAI models, including generative adversarial network, variational autoencoder, transformer and diffusion model [19, 41]. These models find frequent applications in HDT for IoT-healthcare [42, 43, 44, 45].
@@
GAN is significant for HDT in IoT-healthcare, because of its ability to enrich training data, by generating realistic synthetic data. It can ensure patient privacy through anonymized data creation, facilitate data imputation with their cross-modality data translation capabilities, and aid in early anomaly detection for health conditions [42]. For example, a heart dynamics, represented by a ordinary differential equations (ODE) system, was integrated into the training of a GAN to generate biologically plausible electrocardiogram (ECG) data [47], which can be used to solve the ECG data scarcity issue when digital modeling of cardiac activities. Additionally, a chained GAN-based approach, which connected multiple GAN models, was proposed to simulate the pathology of tissues in the human body. Specifically, it included three steps: i) shape generation with a mask generator; ii) initial simulation via domain-specific heuristics; iii) detail refinement with a refining generator. This approach can be used to model tissue digital twin in HDT, as well as help diagnose diseases and predict their progress [48].
@@
VAE is pivotal for HDT in IoT-healthcare. The first key function is compressing and generating complex health data. Additionally, it is also instrumental in anomaly detection, aiding in the early identification of health issues. Moreover, VAE can support health prediction by simulating disease progression [43]. For instance, a VAE-based approach was introduced to reconstruct complete body movements using input signals from the PT’s head-mounted device (HMD) [50]. Specifically, the encoder took the input signals and mapped it into a latent space. Then, this latent space representation was used by the decoder to reconstruct the complete body movements. This method enabled the digital reconstruction of full-body motion based on signals from the head and hands. Similar VAE-based human motion generation was also presented in [51]. These proved beneficial for accurately modeling human motions in scenarios where motion data obtained from the PT may inadequately represent complete body movements, thereby aiding motion monitoring endeavors. Furthermore, a VAE-based method was proposed to predict potential subsequent steps in the clinical measurement trajectory of a patient encountering an ischemic stroke [52]. The model was trained on data from 1216 ischemic stroke patients in the medical information mart for intensive care-IV database. This approach holds promise in simulating disease progression within HDT, thereby assisting in customizing treatment plans tailored to individual patients.
@@
Transformer is well-matched for HDT in IoT-healthcare by adeptly managing sequential health data, such as EHRs and sensor information. The proficiency in sequence-to-sequence tasks allows it to accurately interpret input sequences and generate corresponding outputs. Furthermore, thanks to the multi-head attention mechanism, transformers can masterfully navigate the intricate web of relationships among diverse health data types, e.g., clinical notes and lab results, enhancing their utility in complex healthcare environments [44]. For example, the state-of-the-art language generation model, generative pre-trained transformer (GPT), has been incorporated into EHR workflows to autonomously respond to patients’ healthcare inquiries [54]. This model can be visualized as an HDT representing a doctor, equipped with extensive medical knowledge, capable of addressing patient queries. Additionally, a pre-trained transformer-based approach has been proposed to concurrently learn gene and cell embeddings from a vast amount of single-cell sequencing data. The model further leveraged attention mechanism to capture intricate gene-to-gene interactions at the single-cell level [55]. This approach holds promise in digitally modeling diverse facets of cellular processes within HDT, offering insights into personalized responses to treatments.
@@
Diffusion model is particularly applicable to HDT in IoT-healthcare. It can recreate complex, high-dimensional distributions of data learned from extensive relevant health datasets. This ability enables the diffusion model to capture the intricacies and variations of human health states and processes, thereby generating highly accurate and individual-specific data involved in HDT [45]. For instance, a diffusion model-based approach was proposed to generate individual electroencephalogram (EEG) data [57]. Specifically, the model was trained on maps of electrode-frequency distribution, which were created from extensive EEG datasets labeled with emotions. It can help solve the EEG data scarcity issue in HDT, and enable HDT to digitally model brain activity patterns, supporting the neurological health monitoring. Additionally, by collaborating with an assistive modality embedding as prior information to diffusion model formulation, a diffusion model-based approach was proposed for positron emission tomography (PET) denoising [58]. It can be used to denoise the acquired data with noise, providing cleaned data to HDT for better personalized healthcare monitoring and diagnostic accuracy. Besides, a diffusion-based unconditional 3D human generative model was proposed to 3D human generation [59]. To be more specific, it learned from 2D human images, leveraging the superior modeling and editing capabilities of structured 2D latents over 1D. The method combined a structured 3D-aware auto-decoder with a latent diffusion model to manage the structured 2D latent space effectively. This approach can facilitate the precise digital modeling of 3D human appearances within HDT.
@@ II-C Framework of GAI-driven HDT
Data acquisition component is significantly crucial for HDT, which is the fuel of HDT. The required substantial health data, such as EEG, ECG, and magnetic resonance imaging (MRI) data, for HDT commonly collected from the IoT-enabled pervasive sensing and medical institutions [62]. However, these traditional data acquisitions methods are usually inefficient due to various factors [63, 64]. To this end, GAI can generate ultra-realistic health data based on the collected data for enriching the datasets.
The communication component plays a bridge role in HDT, which is responsible for bi-directional data transmissions between the physical and digital worlds, such as the transmission of the collected data and the feedback in the informal world. However, these data are usually large-scale, multi-modal, and time-sensitive, which are hard to fully met by traditional communication systems [65]. To this end, GAI-enabled communication, such as GAI-enabled semantic communication [66, 67, 68] and cross-modal communication [69], can be applied to support communication in HDT by generating the transmitted data at the received sides, enhancing the data transmission performances.
@@
In the following sections, we will delve into the implementation of GAI-driven HDT in IoT-healthcare.
@@ III GAI-driven HDT Implementation
TABLE IV: Summary of GAI-based Approach for Data Acquisition in HDT.
@@ III-A GAI-enabled Data Acquisition
Data forms the cornerstone of HDT development and service provision [13]. Commonly employed data acquisition methods for HDT primarily stem from diverse sources, including medical institutions [70], as well as both non-invasive and invasive sensors equipped in PTs [40]. Nevertheless, these methods present challenges due to their time-consuming, costly, intrusive nature, and limited scalability. This restricts acquiring extensive individual-level data essential for robust HDT development and comprehensive service delivery. GAI can significantly assist the data acquisition in HDT, by offering diverse and highly realistic synthetic data. In the following, we introduce several common synthetic HDT data generated by GAI, including synthetic physiological, medical imaging and motion data, as summarized in Table IV.
The physiological data, such as ECG, EEG, and photoplethysmogram (PPG), hold significant importance for HDT. By generating synthetic physiological data, GAI mitigates the limitations of traditional physiological data acquisition methods, reducing costs and time associated with collecting extensive real-world physiological data. For instance, SynSigGAN proposed by Hazra et al. in [71] can generate synthesized PPG data, as shown in Fig. 3. In SynSigGAN, the bidirectional grid long short-term memory (LSTM) and the convolutional neural network (CNN) had been used for generator network and discriminator network, respectively, and it was trained on BIDMC PPG and respiration datasets [72]. In addition, diffusion model-based approaches also applied in physiological data generation. Alcaraz et al. in [73] designed a diffusion model-based approach, called structured state space diffusion-ECG (SSSD-ECG), for generating synthetic 12-lead ECG data. SSSD-ECG was built on the SSSDS4 model architecture [74] and trained on PTB-XL dataset [75], which was a publicly available collection of clinical 12-lead ECG data comprising 21,837 records from 18,885 patients. These highly realistic synthetic physiological data can significantly amplify the physiological datasets that required by HDT, to enhance the performance.
@@
The medical imaging data hold significant importance for HDT. Medical imaging provides detailed insights into a PT’s anatomy, allowing for a precise digital representation. This information aids in creating accurate simulations of physiological structures and functions within the VT. Besides, by integrating this data, the VT can simulate and predict the progression of diseases. However, existing medical imaging acquisition methods rely on medical imaging devices, such as MRI or computed tomography (CT) scanners, which encounter high costs, patient discomfort, and safety concerns. To this end, GAI can generate synthetic medical images, augmenting existing datasets without additional patient scans. For instance, Moghadam et al. in [76] proposed a diffusion model-based approach for the synthesis of histopathology images, as shown in Fig. 4. Specifically, the authors used color normalization to force the diffusion model-based approach to learn morphological patterns, and used perception prioritized weighting, aiming to prioritize focusing on diffusion stages with more important structural histopathology contents. Experimental results showed that the proposed approach outperformed the GAN-based approach proposed in [77] by generating high quality histopathology images of brain cancer. While diffusion model and GAN made remarkable progress in medical image generation, they cannot explicitly learn the probability density function of the input data and are highly sensitive to the hyperparameter selections. To mitigate these issues, Hajij et al. in [61] investigated normalizing flows (NFs) based approach as an alternative for synthesizing medical images. Particularly, the authors trained a RealNVP [78], a popular NF model for medical image synthesis, on two medical imaging datasets: chest X-ray [79] and skin cancer [80]. The experimental results showed that the NF-based medical image synthesis approach is an attractive alternative to GAN-based and diffusion model-based approaches.
@@
In summary, GAI offers promising solutions to address the limitations of current data acquisition methods in HDT by generating diverse and high-realistic datasets.